g=9.8

G=6.67×10-11

R=8.40×207

M=?


1
Expert's answer
2021-05-24T12:34:47-0400

Answer

"g=\frac{GM}{r^2}\\9.8=\frac{(6.67\times10^{-11})M}{(8.4\times20^7)2}\\M=1.327\times10^{22}kg"